OpenStudy (anonymous):

PLEASE HELP! This is the only thing keeping from graduating & I never learned at my school What is the distance between points P(0, 4) and Q(3, 0)? Answer 8 7 6 5 What is the distance between P(-3, -2) and Q(5, -2)? Answer 14 12 10 8 Find the midpoint of PQ if P (7.5, 3) and Q(-1.5, 5)? Answer (3, 4) (4, 3) (-4, -3) (-3, 4) Find the midpoint of PQ if P (10, 20) and Q(15, 20)? Answer (2.5, 0) (25, 40) (12.5, 20) (20, 12.5) Find the midpoint of PQ if P (10, 2) and Q(8, 0)? Answer (-9, -1) (-1, 9) (1, 9) (9, 1) Find the midpoint of PQ if P (0, 0) and Q(5, 12)? Answer (2.5, 6) (2, 6) (3.5, 7) (4, 8) . .

OpenStudy (mathmale):

If you'd please choose one problem on which to focus, you'd probably get more responses. After you've chosen one problem, think about what info you'd need to solve that problem. Where would you find that info?

OpenStudy (solomonzelman):

This is definitely not the only thing that keeps you from graduation. I guarantee !!

OpenStudy (anonymous):

distance formula : d = sq rt ((x2 - x1)^2 + (y2 - y1)^2) (0,4) x1 = 0 and y1 = 4 (3,0) x2 = 3 and y2 = 0 now we sub d = sq rt ((3 - 0)^2 + (0 - 4)^2) d = sq rt ((3^2) + (-4^2)) d = sq rt (9 + 16) d = sq rt 25 d = 5 ---- distance between point P and Q is 5 d = sq rt ((x2 - x1)^2 + (y2 - y1)^2) (-3,-2) x1 = -3 and y1 = -2 (5,-2) x2 = 5 and y2 = -2 now we sub d = sq rt ((-2 - (-2)^2 + (5 - (-3)^2) d = sq rt ((-2 + 2)^2 + (5 + 3)^2) d = sq rt (0^2 + 8^2) d = sq rt 64 d = 8 --- distance between P and Q is 8 Midpoint formula : m = (x1 + x2)/2 , (y1 + y2)/2 (7.5,3) x1 = 7.5 and y1 = 3 (-1.5,5) x2 = -1.5 and y2 = 5 now sub m = ((7.5 + (-1.5))/2 , (3 + 5)/2 m = ((7.5 - 1.5)/2 , (8/2) m = 6/2, 8/2 m = (3,4) now that you have the midpoint formula, can you finish the rest ?

OpenStudy (anonymous):

It is the only thing keeping me from graduating, I had to take credit recovery and this is the last test for the last credit I have to make up.

OpenStudy (anonymous):

do you understand ? Its just a matter of subbing in your info into the formula's

OpenStudy (anonymous):

Yes, thank you very much!